traveled 62 miles in 4 hours. some time was biking @ 20mph. the other time spent running @ 8 mph. What is the distance traveled biking? What is the distance traveled biking in the 4 hours if traveled 62 miles

Answer by stanbon(75887)

traveled 62 miles in 4 hours. some time was biking @ 20mph. the other time spent running @ 8 mph. What is the distance traveled biking? What is the distance traveled biking in the 4 hours if traveled 62 miles
----
biking DATA:
distance = x ; rate = 20 mph ; time = x/20 hrs
----
running DATA:
distance = 62-x miles ; rate = 8 mph ; time = (62-x)/8 hrs
--------
Equation:
time + time = 4 hrs
x/20 + (62-x)/8 = 4
Multiply thru by 40 to get:
2x + 5*62-5X = 160
-3x = -150
x = 50 miles (biking distance)
---
62-x = 12 miles (running distance)
=======================================
Cheers,
Stan H.
=============